OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3> -9.1° 1.8° 0.9° 11.8°

OpenStudy (darkbluechocobo):

Question. Is there no picture here?

OpenStudy (anonymous):

no picture

OpenStudy (anonymous):

vector u=-5i+-4j vector v=-4i-3j \[If \theta is the \angle \between the vectors,then\] \[\cos \theta=\frac{ u.v }{ \left| u \right|\left| v \right| }=\frac{ \left( -5i-4j \right).\left( -4i-3j \right) }{\sqrt{\left( 5 \right)^{2}+\left( -4 \right)^{2}} \sqrt{\left( -4\right)^{2} +\left( -3 \right)^{2}} }\] \[\cos \theta=\frac{ \left( -5 \right)\left( -4 \right)+\left( -4 \right)\left( -3 \right) }{ 5\sqrt{41} }\] \[\cos \theta=\frac{ 20+12 }{5\sqrt{41} }=\frac{ 32\sqrt{41} }{ 205 }\] now you can calculate angle by calculator or log tables.

OpenStudy (darkbluechocobo):

Well he has a picture. A very big picture o:

OpenStudy (darkbluechocobo):

Which I cannot help with I haven't taken geometry in 2 years I am sorry ;-;

OpenStudy (anonymous):

so the answer would be 0.9? @surjithayer

OpenStudy (anonymous):

\[if \cos \theta=0.9,then calculate \cos^{-1} \left( 0.9 \right)\]

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

yw